# Private Continual Counting of Unbounded Streams

### v1

@@ Abstract
We study the problem of differentially private continual counting in the unbounded setting where the input size $n$ is not known in advance. Current state-of-the-art algorithms based on optimal instantiations of the matrix mechanism [33] cannot be directly applied here because their privacy guarantees only hold when key parameters are tuned to $n$. Using the common ‘doubling trick’ avoids knowledge of $n$ but leads to suboptimal and non-smooth error. We solve this problem by introducing novel matrix factorizations based on logarithmic perturbations of the function $\frac{1}{\sqrt{1-z}}$ studied in prior works, which may be of independent interest. The resulting algorithm has smooth error, and for any $\alpha>0$ and $t\leq n$ it is able to privately estimate the sum of the first $t$ data points with $O(\log^{2+2\alpha}(t))$ variance. It requires $O(t)$ space and amortized $O(\log t)$ time per round, compared to $O(\log(n)\log(t))$ variance, $O(n)$ space and $O(n\log n)$ pre-processing time for the nearly-optimal bounded-input algorithm of Henzinger et al. [27]. Empirically, we find that our algorithm’s performance is also comparable to theirs in absolute terms: our variance is less than 1.5×1.5$\times$1.5 × theirs for $t$ as large as $2^{24}$.
@@ 1 Introduction
Differentially private counting under continual observation [11, 18] refers to the problem of maintaining accurate running totals over streams of sensitive data. It has attracted a great deal of attention in recent years [3, 4, 6, 7, 14, 16, 23, 27, 28, 35] for its wide-ranging applications in optimization [13, 12, 14, 21, 32, 36], as well as online learning and the private analysis of streaming data more broadly [1, 10, 30, 38, 40]. It has also been used in many large-scale deployments of differential privacy, such as Google’s private next-word prediction [34] and Smart Select models [26]. To a large extent, this flurry of activity has been prompted by recent algorithmic improvements in matrix factorizations for streaming data [14, 23, 27, 33], which have dramatically improved privacy/utility tradeoffs compared to classical approaches based on binary trees [11, 18, 29].
@@
One challenge when applying matrix-based algorithms to streaming settings is that they largely assume access to the input size $n$ as a parameter. While this is a very natural assumption in contexts like private model training, researchers from the beginning have also motivated their work with appeals to applications like public health [18] where it is much less clear how the assumption could hold. In fact, it is known that difficulties related to unbounded inputs have influenced practical deployments of DP [2]: famously, after Apple announced that it planned to use differential privacy to collect user data, Tang et al. [39] discovered that their implementation achieved this by guaranteeing privacy only over a single day while allowing privacy loss per-user to grow without bound over time.
@@
Simple solutions to this problem all carry fundamental tradeoffs. Lifting a bounded-input algorithm to the unbounded case with some form of doubling trick [11] asymptotically preserves performance, but it also leads to non-smooth growth in error over time, undermining one of the major selling points of matrix methods over binary trees [27]. Using a bounded matrix mechanism alongside the classical query model [17, 19] and refusing to produce additional outputs once our privacy budget is exhausted preserves performance and smoothness, but is not truly unbounded and is unlikely to be satisfying in practice. Finally, simple algorithms like adding independent noise to each data point are trivially unbounded with smooth noise, but have extremely sub-optimal error. Our main contribution in this paper is a matrix-based algorithm that overcomes this trilemma: to the best of our knowledge, it is the first private counting algorithm that is simultaneously smooth and unbounded with almost-optimal asymptotic error. It is also practical to implement, with essentially the same complexity as the foundational algorithm of Henzinger et al. [27], and we show empirically that it enjoys small error for realistic input sizes and not just asymptotically.
@@ 2.1 Background
In continual counting [18, 11], the input consists of a (potentially unbounded) stream of data $x_{1},x_{2},\ldots$. At each time step $t$, the algorithm receives $x_{t}$ and then produces an estimate of the partial sum $S_{t}=\sum_{i=1}^{t}x_{t}$. For the sake of simplicity, we will assume that $x_{t}$ is a single bit, but our results are easy to generalize to $x_{t}\in\mathbb{R}^{d}$.
@@
Define two (potentially unbounded) streams $x_{1},x_{2},\ldots$ and $x^{\prime}_{1},x^{\prime}_{2},\ldots$ to be neighboring if there is at most one time step $t$ where $x_{t}\neq x^{\prime}_{t}$. Let $X_{n}$ (resp. $X^{\prime}_{n}$) denote the first $n$ data points in each stream. We say that a mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy (DP) [20] in the continual observation model [18] if for all neighboring streams $X,X^{\prime}$, all $n\in\mathbb{Z}^{+}$, and all measurable events $S$, we have $\mathbb{P}(\mathcal{M}(X_{n})\in S)\leq e^{\varepsilon}\mathbb{P}(\mathcal{M}(X^{\prime}_{n}))+\delta$. Here, $\mathcal{M}(X)$ represents the full sequence of outputs produced by $\mathcal{M}$ on input $X$. This reduces to the standard definition when the streams are finite; the additional quantification over $n$ is necessary to make sense of the definition for unbounded streams. The mechanisms we study all use additive Gaussian noise, and so can readily be shown to satisfy other notions of indistinguishability as well, such as Rényi DP [37], zCDP [9], or Gaussian DP [15]. By virtue of using Gaussian noise, they also automatically satisfy DP in the stronger adaptive setting where $x_{t}$ is permitted to depend on earlier outputs, which is important for learning applications  [14].
Two measures of utility are standard in the literature: the expected mean-squared additive error and the expected maximum additive error, defined respectively as the maximums over $X_{n}$ of:
$\operatorname{err}^{\ell_{2}^{2}}_{n}(\mathcal{M}):=\frac{1}{n}\mathbb{E}_{\mathcal{M}}\big{(}\sum_{t\leq n}(\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{2}^{2}\big{)}^{1/2}$ $\operatorname{err}^{\ell^{\infty}}_{n}(\mathcal{M}):=\mathbb{E}_{\mathcal{M}}\big{(}\max_{t\leq n}~{}\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{\infty}\big{)}$
When $\mathcal{M}$ uses additive Gaussian noise, both metrics can be easily derived from $\mathbb{V}(\mathcal{M}(X_{t}))$, the variance of the noise added as a function of time. This is the main metric we will report when comparing the performance of different algorithms.
@@ 2.2 Related Works
Differential privacy under continual observation was introduced by Dwork et al. [18] and Chan et al. [11], who independently proposed the binary mechanism. Conceptually, this algorithm computes partial sums with the help of a binary tree where leaf nodes contain private estimates of individual data points, and internal nodes contain independent estimates of the subsequences spanned by their children. The binary mechanism of Dwork et al. [18] used Laplacian noise to satisfy $(\varepsilon,0)$-DP with $\operatorname{err}_{n}^{\ell^{\infty}}=O(\log^{2.5}(n))$, which was improved to $O(\log^{2.5}(t))$ by Chan et al. [11]. This was later reduced to $O(\log(t)\sqrt{\log n})$ by Jain et al. [30] by using Gaussian noise instead. A straightforward optimization of the algorithm follows from the observation that at a given time step, it is only necessary to store $\log_{2}(n)$ nodes of the tree in memory. Subsequent work by Honaker [29] additionally showed that it is possible to reduce the error by roughly a factor of 2 by making efficient use of multiple independent estimates for the same partial sum.
@@
Both the original binary mechanism and Honaker’s variant have very uneven noise scale across time steps. Recently, Andersson and Pagh [3] proposed a smooth variant of the binary mechanism which equalizes the noise distribution across time steps, further improving $\operatorname{err}^{\ell_{\infty}}_{n}$ by a constant factor.
@@
The binary mechanism can be seen as a special case of the general matrix method  [33]. Given a dataset $x$ and a matrix $A=LR$, the matrix method estimates $Ax$ by $L(Rx+y)=Ax+Ly$. This satisfies DP when $y$ is a Gaussian scaled to $\rVert R\rVert_{1\to 2}$, the maximum $\ell_{2}$ norm over the columns of $R$. The final error also scales with $\lVert L\rVert_{2\to\infty}$, the maximum $\ell_{2}$ norm over the rows of $L$, and so by choosing $L$ and $R$ carefully we can achieve error depending on the factorization norm $\gamma_{2}(A)=\min\{\lVert L\rVert_{2\to\infty}\lVert R\rVert_{1\to 2}\mid LR=A\}$. This is often much more accurate than calibrating $y$ to $A$ directly. In the case of continual counting, $A=M_{\textit{count}}$, the all-ones lower-triangular matrix.
@@
This matrix-based approach has rapidly become the standard in the field. Edmonds et al. [22] proved a lower bound of $\Omega(\log n)$ on the per-time-step $\ell_{\infty}$ error of continual counting by lower-bounding $\gamma_{2}(M_{\textit{count}})$. Denisov et al. [14] approached the problem of finding optimal decompositions as a convex optimization problem, building on earlier work in the offline setting  [41]. They proved that, for lower-triangular matrices like $M_{\textit{count}}$, it is sufficient to consider only decompositions where $L$ and $R$ are both lower-triangular. This is a particularly nice result because it implies there is no extra cost for using the matrix method in streaming settings. Subsequently, Fichtenberger et al. [23] provided an explicit factorization $L=R=M_{\textit{count}}^{1/2}$ and studied its $\operatorname{err}^{\ell_{\infty}}_{n}$ error, and Henzinger et al. [27] proved that the same decomposition is simultaneously nearly optimal in terms of $\operatorname{err}^{\ell_{2}^{2}}_{n}$. More recent work has investigated various computational modifications of this factorization, trading a small amount of accuracy for efficiency by approximating the optimal decomposition in more compact ways, e.g. through finite recurrences [16, 5] or by binning together similar values [4].
@@
A distinct line of work has investigated the unbounded case. In their seminal work, Chan et al. [11] proposed a ‘hybrid’ mechanism which uses a variant of the classic doubling trick to match the asymptotic $\operatorname{err}_{n}^{\ell_{\infty}}$ rate of the binary mechanism when $n$ is unknown. Their technique is generic and can be used to ‘lift’ any fixed-size algorithm into one that works on unbounded streams with the help of a (potentially inefficient) unbounded algorithm — in section 5 we investigate a version of this mechanism instantiated with Algorithm 1. An alternative strategy for handling unbounded inputs is to only permit queries based on sliding windows or decaying sums [8, 40]. The dual approach, proposed by Bolot et al. [8] and recently extended by Andersson et al. [6], instead relaxes the classical privacy definition to allow DP guarantees for past data points to gradually expire over time.
@@
Our work is most directly inspired by the rational approximation methods of Dvijotham et al. [16], and in particular by their use of the equivalence between certain matrix factorizations and generating functions. While they search for simple approximations of functions to derive computationally efficient factorizations, we take the opposite approach by searching for more complicated functions to derive matrix factorizations that are usable in the unbounded setting. Because we are not directly concerned with reducing computational complexity below $O(n)$ space and $O(n\log n)$ time, we will generally take the $M_{\textit{count}}^{1/2}$ mechanism of Henzinger et al. [27] and Fichtenberger et al. [23] (“Sqrt Matrix” in Table 1 and Figure 2) as the baseline against which our algorithm is compared.
@@
Mechanism $\mathbb{V}(\mathcal{M}(X_{t}))$ Time for $t$ outputs Space Smooth Unbounded
Binary  [11, 30]* $O(\log(t)\log(n))$ $O(t)$ $O(\log t)$ No No
Hybrid Binary  [11]* $O(\log^{2}t+\log t)$ $O(t)$ $O(\log t)$ No Yes
Smooth Binary [3] $O((\log n+\log\log n)^{2})$ $O(t)$ $O(\log t)$ Yes No
Sqrt Matrix  [27, 23] $O(\log(t)\log(n))$ $O(n\log n)$ $O(n)$ Yes No

### v2

@@ Abstract
We study the problem of differentially private continual counting in the unbounded setting where the input size $n$ is not known in advance. Current state-of-the-art algorithms based on optimal instantiations of the matrix mechanism [li_matrix_nodate] cannot be directly applied here because their privacy guarantees only hold when key parameters are tuned to $n$. Using the common ‘doubling trick’ avoids knowledge of $n$ but leads to suboptimal and non-smooth error. We solve this problem by introducing novel matrix factorizations based on logarithmic perturbations of the function $\frac{1}{\sqrt{1-z}}$ studied in prior works, which may be of independent interest. The resulting algorithm has smooth error, and for any $\alpha>0$ and $t\leq n$ it is able to privately estimate the sum of the first $t$ data points with $O(\log^{2+2\alpha}(t))$ variance. It requires $O(t)$ space and amortized $O(\log t)$ time per round, compared to $O(\log(n)\log(t))$ variance, $O(n)$ space and $O(n\log n)$ pre-processing time for the nearly-optimal bounded-input algorithm of henzinger_almost_2023. Empirically, we find that our algorithm’s performance is also comparable to theirs in absolute terms: our variance is less than $1.5\times$ theirs for $t$ as large as $2^{24}$.
@@ 1 Introduction
Differentially private counting under continual observation [chan2011private, dwork_differential_nodate] refers to the problem of maintaining accurate running totals over streams of sensitive data. It has attracted a great deal of attention in recent years [andersson_smooth_nodate, andersson_streaming_2024, andersson_continual_nodate, bienstock_dmm_2024, denisov2022improved, dvijotham_efficient_2024, fichtenberger2023constant, henzinger_almost_2023, henzinger2024unifying, mcmahan_inversion_2025] for its wide-ranging applications in optimization [choquette2023amplified, choquette2022multi, denisov2022improved, dwork2014analyze, kairouz2021practical, mcmahan2024hassle], as well as online learning and the private analysis of streaming data more broadly [agarwal2017price, cardoso2022differentially, jain_differentially_nodate, smith_nearly_nodate, upadhyay2019sublinear]. It has also been used in many large-scale deployments of differential privacy, such as Google’s private next-word prediction [mcmahan2022federated] and Smart Select models [hartmann2023distributed]. To a large extent, this flurry of activity has been prompted by recent algorithmic improvements in matrix factorizations for streaming data [denisov2022improved, fichtenberger2023constant, henzinger_almost_2023, li_matrix_nodate], which have dramatically improved privacy/utility tradeoffs compared to classical approaches based on binary trees [chan2011private, dwork_differential_nodate, honaker2015efficient].
@@
One challenge when applying matrix-based algorithms to streaming settings is that they largely assume access to the input size $n$ as a parameter. While this is a very natural assumption in contexts like private model training, researchers from the beginning have also motivated their work with appeals to applications like public health [dwork_differential_nodate] where it is much less clear how the assumption could hold. In fact, it is known that difficulties related to unbounded inputs have influenced practical deployments of DP [amin2024practical]: famously, after Apple announced that it planned to use differential privacy to collect user data, tang2017privacy discovered that their implementation achieved this by guaranteeing privacy only over a single day while allowing privacy loss per-user to grow without bound over time.
@@
Simple solutions to this problem all carry fundamental tradeoffs. Lifting a bounded-input algorithm to the unbounded case with some form of doubling trick [chan2011private] asymptotically preserves performance, but it also leads to non-smooth growth in error over time, undermining one of the major selling points of matrix methods over binary trees [henzinger_almost_2023]. Using a bounded matrix mechanism alongside the classical query model [dwork2004privacy, dwork2006calibrating] and refusing to produce additional outputs once our privacy budget is exhausted preserves performance and smoothness, but is not truly unbounded and is unlikely to be satisfying in practice. Finally, simple algorithms like adding independent noise to each data point are trivially unbounded with smooth noise, but have extremely sub-optimal error. Our main contribution in this paper is a matrix-based algorithm that overcomes this trilemma: to the best of our knowledge, it is the first private counting algorithm that is simultaneously smooth and unbounded with almost-optimal asymptotic error. It is also practical to implement, with essentially the same complexity as the foundational algorithm of henzinger_almost_2023, and we show empirically that it enjoys small error for realistic input sizes and not just asymptotically.
@@ 2.1 Background
In continual counting [dwork_differential_nodate, chan2011private], the input consists of a (potentially unbounded) stream of data $x_{1},x_{2},\ldots$. At each time step $t$, the algorithm receives $x_{t}$ and then produces an estimate of the partial sum $S_{t}=\sum_{i=1}^{t}x_{t}$. For the sake of simplicity, we will assume that $x_{t}$ is a single bit, but our results are easy to generalize to $x_{t}\in\mathbb{R}^{d}$.
@@
Define two (potentially unbounded) streams $x_{1},x_{2},\ldots$ and $x^{\prime}_{1},x^{\prime}_{2},\ldots$ to be neighboring if there is at most one time step $t$ where $x_{t}\neq x^{\prime}_{t}$. Let $X_{n}$ (resp. $X^{\prime}_{n}$) denote the first $n$ data points in each stream. We say that a mechanism $\mathcal{M}$ satisfies $(\varepsilon,\delta_{priv})$-differential privacy (DP) [dwork2014algorithmic] in the continual observation model [dwork_differential_nodate] if for all neighboring streams $X,X^{\prime}$, all $n\in\mathbb{Z}^{+}$, and all measurable events $S$, we have $\mathbb{P}(\mathcal{M}(X_{n})\in S)\leq e^{\varepsilon}\mathbb{P}(\mathcal{M}(X^{\prime}_{n}))+\delta_{priv}$. Here, $\mathcal{M}(X)$ represents the full sequence of outputs produced by $\mathcal{M}$ on input $X$. This reduces to the standard definition when the streams are finite; the additional quantification over $n$ is necessary to make sense of the definition for unbounded streams. The mechanisms we study all use additive Gaussian noise, and so can readily be shown to satisfy other notions of indistinguishability as well, such as Rényi DP mironov2017renyi, zCDP bun2016concentrated, or Gaussian DP dong2022gaussian. By virtue of using Gaussian noise, they also automatically satisfy DP in the stronger adaptive setting where $x_{t}$ is permitted to depend on earlier outputs, which is important for learning applications  [denisov2022improved].
Two measures of utility are standard in the literature: the root mean squared additive error and the expected maximum additive error, defined respectively as the maximums over $X_{n}$ of:
$\operatorname{err}^{\ell_{2}}_{n}(\mathcal{M}):=\sqrt{\frac{1}{n}\mathbb{E}_{\mathcal{M}}\big(\sum_{t\leq n}(\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{2}^{2}\big)}$ $\operatorname{err}^{\ell_{\infty}}_{n}(\mathcal{M}):=\mathbb{E}_{\mathcal{M}}\big(\max_{t\leq n}~\lVert S_{t}-\mathcal{M}(X_{t})\rVert_{\infty}\big)$
When $\mathcal{M}$ uses additive Gaussian noise, both expectations and matching high-probability bounds can be derived from $\mathbb{V}(\mathcal{M}(X_{t}))$, the (non-asymptotic) variance of the noise added as a function of time. This is the main metric we will report when comparing the performance of different algorithms. We additionally say that a mechanism has smooth error if $\mathbb{V}(\mathcal{M}(X_{t}))$ is a smooth function of $t$. This is a desirable property because $\operatorname{err}^{\ell_{\infty}}_{n}$ depends on the maximum variance across all time steps, which can be significantly higher than the average variance when error is non-smooth [andersson_smooth_nodate].
@@ 2.2 Related Works
Differential privacy under continual observation was introduced by dwork_differential_nodate and chan2011private, who independently proposed the binary mechanism. Conceptually, this algorithm computes partial sums with the help of a binary tree where leaf nodes contain private estimates of individual data points, and internal nodes contain independent estimates of the subsequences spanned by their children. The binary mechanism of dwork_differential_nodate used Laplacian noise to satisfy $(\varepsilon,0)$-DP with $\operatorname{err}_{n}^{\ell^{\infty}}=O(\log^{2.5}(n))$, which was improved to $O(\log^{2.5}(t))$ by chan2011private. This was later reduced to $O(\log(t)\sqrt{\log n})$ by jain_differentially_nodate by using Gaussian noise instead. A straightforward optimization of the algorithm follows from the observation that at a given time step, it is only necessary to store $\log_{2}(n)$ nodes of the tree in memory. Subsequent work by honaker2015efficient additionally showed that it is possible to reduce the error by roughly a factor of 2 by making efficient use of multiple independent estimates for the same partial sum.
@@
Both the original binary mechanism and Honaker’s variant have very uneven noise scale across time steps. Recently, andersson_smooth_nodate proposed a smooth variant of the binary mechanism which equalizes the noise distribution across time steps, further improving $\operatorname{err}^{\ell_{\infty}}_{n}$ by a constant factor.
@@
The binary mechanism can be seen as a special case of the general matrix method  [li_matrix_nodate]. Given a dataset $x$ and a matrix $A=LR$, the matrix method estimates $Ax$ by $L(Rx+z)=Ax+Lz$. This satisfies DP when $z$ is a Gaussian scaled to $\rVert R\rVert_{1\to 2}$, the maximum $\ell_{2}$ norm over the columns of $R$. The final error also scales with $\lVert L\rVert_{2\to\infty}$, the maximum $\ell_{2}$ norm over the rows of $L$, and so by choosing $L$ and $R$ carefully we can achieve error depending on the factorization norm $\gamma_{2}(A)=\min\{\lVert L\rVert_{2\to\infty}\lVert R\rVert_{1\to 2}\mid LR=A\}$. This is often much more accurate than calibrating $z$ to $A$ directly. In the case of continual counting, $A=M_{\textit{count}}$, the all-ones lower-triangular matrix.
@@
This matrix-based approach has rapidly become the standard in the field. edmonds_power_2019 proved a lower bound of $\Omega(\log n)$ on the per-time-step $\ell_{\infty}$ error of continual counting by lower-bounding $\gamma_{2}(M_{\textit{count}})$. denisov2022improved approached the problem of finding optimal decompositions as a convex optimization problem, building on earlier work in the offline setting  [yuan2016optimal]. They proved that, for lower-triangular matrices like $M_{\textit{count}}$, it is sufficient to consider only decompositions where $L$ and $R$ are both lower-triangular. This is a particularly nice result because it implies there is no extra cost for using the matrix method in streaming settings. Subsequently, fichtenberger2023constant provided an explicit factorization $L=R=M_{\textit{count}}^{1/2}$ and studied its $\operatorname{err}^{\ell_{\infty}}_{n}$ error, and henzinger_almost_2023 proved that the same decomposition is simultaneously nearly optimal in terms of $\operatorname{err}^{\ell_{2}}_{n}$. More recent work has investigated various computational modifications of this factorization, trading a small amount of accuracy for efficiency by approximating the optimal decomposition in more compact ways, e.g. through finite recurrences [dvijotham_efficient_2024, andersson_space_2025] or by binning together similar values [andersson_streaming_2024].
@@
A distinct line of work has investigated the unbounded case. In their seminal work, chan2011private proposed a ‘hybrid’ mechanism which uses a variant of the classic doubling trick to match the asymptotic $\operatorname{err}_{n}^{\ell_{\infty}}$ rate of the binary mechanism when $n$ is unknown. Their technique is generic and can be used to ‘lift’ any fixed-size algorithm into one that works on unbounded streams with the help of a (potentially inefficient) unbounded algorithm — in section 5 we investigate a version of this mechanism instantiated with Algorithm 1. An alternative strategy for handling unbounded inputs is to only permit queries based on sliding windows or decaying sums [bolot2013private, upadhyay2019sublinear]. The dual approach, proposed by bolot2013private and recently extended by andersson_continual_nodate, instead relaxes the classical privacy definition to allow DP guarantees for past data points to gradually expire over time.
@@
Our work is most directly inspired by the rational approximation methods of dvijotham_efficient_2024, and in particular by their use of the equivalence between certain matrix factorizations and generating functions. While they search for simple approximations of functions to derive computationally efficient factorizations, we take the opposite approach by searching for more complicated functions to derive matrix factorizations that are usable in the unbounded setting. Because we are not directly concerned with reducing computational complexity below $O(n)$ space and $O(n\log n)$ time, we will generally take the $M_{\textit{count}}^{1/2}$ mechanism of henzinger_almost_2023 and fichtenberger2023constant (“Sqrt Matrix” in Table 1 and Figure 2) as the baseline against which our algorithm is compared.
@@
Mechanism $\mathbb{V}(\mathcal{M}(X_{t}))\cdot\frac{\varepsilon^{2}}{\log(1/\delta_{priv})}$ Time for $t$ outputs Space Smooth Unbounded
Binary  [chan2011private, jain_differentially_nodate]* $O(\log(t)\log(n))$ $O(t)$ $O(\log t)$ No No
Hybrid Binary  [chan2011private]* $O(\log^{2}t+\log t)$ $O(t)$ $O(\log t)$ No Yes
Smooth Binary [andersson_smooth_nodate] $O((\log n+\log\log n)^{2})$ $O(t)$ $O(\log t)$ Yes No
Sqrt Matrix  [henzinger_almost_2023, fichtenberger2023constant] $O(\log(t)\log(n))$ $O(n\log n)$ $O(n)$ Yes No
Algorithm 1 $O(\log^{2+2\alpha}t)$ $O(t\log t)$ $O(t)$ Yes Yes